# Siegel modular forms arising from higher Chow cycles

### v1

@@ 1. Introduction
One might think, like myself when started this project, that the same construction is possible for other types of cycles and arbitrary $g$ to produce plenty Siegel modular forms. Indeed, we have the infinitesimal invariant for any type of higher cycles. However, it turns out from the calculation of Nori ([24] p.372) that the Koszul cohomology sheaf where the primitive infinitesimal invariant lives is nonvanishing only in the above cases, beyond the classical case of 00-cycles (and divisors). I now understand that this is a sort of exceptional phenomenon in low dimension, like the surjectivity of the Torelli map in $g\leq 3$.
Theorem 1.1 is not surprising if we view the Koszul complex as a complex of automorphic vector bundles. The boundary behavior of $f_{Z}$ is a consequence of the admissibility of the normal function of $Z$ ([13], [26]). Nevertheless, in view of the current distance of the two branches, it would not be entirely trivial to make this link explicit, which will serve as the basis of further investigation. An observation similar to Theorem 1.1 (but more primitive) was made for $K3$ surfaces in [20] §3.8.
Our more practical effort in this paper is devoted to showing that the correspondence $Z\rightsquigarrow f_{Z}$ is functorial with respect to degeneration. Let $g\geq 2$. We choose a corank $1$ (i.e., maximal) cusp ${\mathcal{A}_{I}}$ of ${\mathcal{A}_{\Gamma}}$, which itself is a Siegel modular variety of genus $g-1$. In the modular form side, we have the Siegel operator $\Phi_{I}$ which is a sort of restriction operator to ${\mathcal{A}_{I}}$ (see [28], [12]). This produces vector-valued Siegel modular forms on ${\mathcal{A}_{I}}$ from those on ${\mathcal{A}_{\Gamma}}$. We extend this operation to meromorphic modular forms (§5). On the other hand, we have the rank $1$ degeneration of the universal abelian variety over the toroidal boundary divisor over ${\mathcal{A}_{I}}$. For a cycle-theoretic technical reason, we restrict ourselves to the case of irreducible degeneration (cf. Remark 6.3). Then, for a cycle family $Z$ as above, we have a degeneration operation known as the K-theory elevator ([7], [10]). In an (over)simplified setting, this produces a family $Z_{I}$ of higher Chow cycles on the universal family over ${\mathcal{A}_{I}}$ whose K-theoretic degree increases by $1$, namely if $Z$ has type $(2,n)$ then $Z_{I}$ has type $(2,n+1)$.
@@
The proof of Theorem 6.2 is done by comparing the Siegel operator with the limit formula of [10] for the normal functions. The partial toroidal compactification serves as a common place for both operations. It turns out that the reduction of automorphic vector bundles which takes place in the process of Siegel operator agrees with the assumption made in [10] on the Kunneth type of differential forms.
@@
    Siegel operator for meromorphic modular forms (§5)
@@
The motivation of this paper comes from the pioneering work of Collino and Pirolla ([7], [8]). In [8], they discovered that the infinitesimal invariant of the Ceresa cycle in genus $3$ is essentially the equation of the universal plane quartic. Much later, the latter was studied independently from the viewpoint of modular forms ([6]). In the subsequent paper [7], Collino constructed higher Chow cycles in $g=2$ and $g=1$ by applying the K-theory elevator to the Ceresa cycle in $g=3$ successively, and expressed their infinitesimal invariants in terms of invariants of six points on ${{\mathbb{P}}}^{1}$ in the case $g=2$, and theta series in the case $g=1$ respectively. It was observed, again later and independently, that invariants of binary sextics are vector-valued Siegel modular forms ([5]). In a sense, like the age of Lemniscate, algebraic cycles have been ahead of modular forms. The present paper grew out of an effort to understand the idea of Collino from the viewpoint of modular forms.
@@ 3. Higher Chow cycles
In this section we recall higher Chow cycles and normal functions. Our main references are [19] and [13]. Note that we would not need to define higher Chow groups, and hence avoid technicalities regarding the moving lemma in our presentation. We use the cubical presentation of higher Chow cycles.
@@ 3.1. Higher Chow cycles
For nullhomologous $Z$, the regulator $\nu(Z)\in J^{p,n}(X)$ is also called the Abel-Jacobi invariant of $Z$. This is a generalization of the classical Abel-Jacobi invariant in the case $n=0$. An integral formula in the case $n=1$ has been classically known as Levine’s formula; a generalization to the case $n\geq 2$ was given by Kerr, Lewis and Müller-Stach [19]. In this formula, a further proper intersection condition is imposed on $Z$ with regard to the real faces of ${\square}^{n}$ (see [19] §5.4). In §6 and later, where we implicitly use this formula, we tacitly assume that our cycles satisfy this condition.
@@ 4. The main construction
In this section we present and prove the full version of Theorem 1.1. In §4.1 we calculate the relevant Koszul complex as a complex of automorphic vector bundles. Our main construction is given in §4.2.
@@ 4.2. The main construction
We finish this section with some comments on other possible constructions.
@@
When $g\geq 2$, we could also look at the non-primitive part of $\delta\nu_{Z}$. By (4.5), this produces Siegel modular forms of (virtual) weight $\wedge^{2}{{\rm Sym}}^{2}\otimes\det^{-1}$ when $g=3$, and weight ${{\rm Sym}}^{2}$ when $g=2$.
@@
When the etale cover $S$ itself has a nice moduli interpretation and a satisfactory theory of “modular forms” on $S$ is available, we could directly study $\delta\nu_{Z}$ itself as such a generalized modular form on $S$ without pushing it down to ${\mathcal{A}_{\Gamma}}$. This is not the approach taken in this paper, but what we have in mind is the moduli space $\mathcal{M}_{3}\to\mathcal{A}_{3}$ of genus $3$ curves, for which the theory of Teichmüller modular forms is known ([18], [6]).
@@
Another interesting source of normal functions is the extension classes of the mixed Hodge structures on the fundamental groups of curves ([15]). It is known that the first nontrivial extension class corresponds to the Ceresa cycle in the case $g=3$ ([25]) and to the Collino cycle in the case $g=2$ ([9]). Probably the higher extension classes also give Siegel modular forms, but I do not know whether the admissibility is known for them.
@@ 6.1. Rank $1$ degeneration
We will come back to (1) – (5) from a general point of view on another occasion. For the moment, these are taken as assumptions in the rest of this paper.
@@ 6.3. The statement
Recall from Theorem 4.2 that for each $1\leq i\leq d$, the cycle family $Z$ gives a meromorphic Siegel modular form $f_{Z}^{(i)}$ of weight ${{\rm Sym}}^{4i}\otimes{\rm det^{-i}}$ on ${\mathcal{A}_{\Gamma}}$. In Lemma 8.1, we will prove that $f_{Z}^{(i)}$ is holomorphic at a general point of $\Delta_{I}$. However, as explained in §5.3, this is still weaker than the regularity at $I$ in the sense of Definition 5.4. Let us assume that the divisor components of ${\mathcal{A}_{\Gamma}}-U$ do not contain ${\mathcal{A}_{I}}$ in their closure. Then, by Lemma 5.3 (1), $f_{Z}^{(i)}$ is regular at $I$. This enables us to apply the meromorphic Siegel operator to $f_{Z}^{(i)}$. By Proposition 5.6, the resulting $\Phi_{I}f_{Z}^{(i)}$ is a meromorphic Siegel modular form over ${\mathcal{A}_{I}}$, again of weight ${{\rm Sym}}^{4i}\otimes{\rm det}^{-i}$.
On the other hand, if consider the cycle family $Z_{I}$ on $X_{I}\to S_{I}$ and take the same product construction as $\delta\nu_{Z}^{+}\rightsquigarrow f_{Z}^{(i)}$ for the etale cover $S_{I}\to{\widetilde{\mathcal{A}}_{I}}$ of degree $d$ (rather than $S_{I}\to{\mathcal{A}_{I}}$ which in general has larger degree), we obtain a meromorphic section $f_{Z_{I}}^{(i)}$ of $\phi_{0}^{\ast}({{\rm Sym}}^{4i}{\mathcal{E}}_{I}\otimes{\mathcal{L}}_{I}^{\otimes-i})$ over ${\widetilde{\mathcal{A}}_{I}}$, where ${\mathcal{E}}_{I}$, ${\mathcal{L}}_{I}$ are the Hodge bundles on ${\mathcal{A}_{I}}$.
@@ 8.4. The case $g=3$
The proof of Theorem 6.2 in the case $g=3$ is similar to the case $g=2$. In this case $\Phi_{I}H^{1}\mathcal{K}_{3}$ no longer coincides with $\Phi_{I}\mathcal{H}^{1,2}_{{\rm prim}}\otimes\Phi_{I}\Omega_{{\mathcal{D}}}^{1}$, but is still realized as a sub vector bundle of $\Phi_{I}\mathcal{H}^{1,2}_{{\rm prim}}\otimes\Phi_{I}\Omega_{{\mathcal{D}}}^{1}$ by Proposition 7.4. The cusp ${\mathcal{D}}_{I}$ is now the $3$-dimensional Siegel upper half space, so we use $3$-variable coordinates $(\tau_{ij})_{i,j}$ instead of a single variable $\tau$. Then, replacing $\partial_{\tau}\otimes\omega$ by $\partial_{\tau_{ij}}\otimes\omega$ as the test dual sections where now $\omega_{t}\in H^{2,1}_{{\rm prim}}(X_{t})$, the proof of Proposition 8.4 is valid in the case $g=3$. The step deducing Theorem 6.2 from Proposition 8.4 remains the same.

### v2

@@ 1. Introduction
One might think, like myself when started this project, that the same construction is possible for other types of cycles and arbitrary $g$ to produce plenty Siegel modular forms. Indeed, we have the infinitesimal invariant for any type of cycles. However, it turns out from the calculation of Nori ([28] p.372) that the Koszul cohomology sheaf where the primitive infinitesimal invariant lives is nonvanishing only in the above cases, beyond the classical case of 00-cycles (and divisors). I now understand that this is a sort of exceptional phenomenon in low dimension, like the surjectivity of the Torelli map in $g\leq 3$. Moreover, by some rigidity theorems in the theory of algebraic cycles, we find that only countably many modular forms can be obtained in this way (§4.4). Thus these modular forms should be precious.
Theorem 1.1 is not surprising if we view the Koszul complex as a complex of automorphic vector bundles. The boundary behavior of $f_{Z}$ is a consequence of the admissibility of the normal function of $Z$ ([14], [30]). Nevertheless, in view of the current distance of the two branches, it would not be entirely trivial to make this link explicit, which will serve as the basis of further investigation. In the case $g=3$, a similar observation was also made recently by Hain [18] in the context of Teichmüller modular forms. In a more primitive form, this type of observation was done for $K3$ surfaces in [23] §3.8.
Our more practical effort in this paper is devoted to showing that the correspondence $Z\rightsquigarrow f_{Z}$ is functorial with respect to degeneration. Let $g\geq 2$. We choose a corank $1$ (i.e., maximal) cusp ${\mathcal{A}_{I}}$ of ${\mathcal{A}_{\Gamma}}$, which itself is a Siegel modular variety of genus $g-1$. In the modular form side, we have the Siegel operator $\Phi_{I}$ which is a sort of restriction operator to ${\mathcal{A}_{I}}$ (see [32], [13]). This produces vector-valued Siegel modular forms on ${\mathcal{A}_{I}}$ from those on ${\mathcal{A}_{\Gamma}}$. We extend this operation to meromorphic modular forms (§5). On the other hand, we have the rank $1$ degeneration of the universal abelian variety over the toroidal boundary divisor over ${\mathcal{A}_{I}}$. For a cycle-theoretic technical reason, we restrict ourselves to the case of irreducible degeneration (cf. Remark 6.3). Then, for a cycle family $Z$ as above, we have a degeneration operation known as the K-theory elevator ([8], [11]). In an (over)simplified setting, this produces a family $Z_{I}$ of higher Chow cycles on the $(g-1)$-dimensional universal family over ${\mathcal{A}_{I}}$ whose K-theoretic degree increases by $1$, namely if $Z$ has type $(2,n)$ then $Z_{I}$ has type $(2,n+1)$.
@@
The proof of Theorem 6.2 is done by comparing the Siegel operator with the limit formula of [11] for the normal functions. The partial toroidal compactification serves as a common place for both operations. A point of the proof is that the reduction of automorphic vector bundles which takes place in the process of Siegel operator agrees with the assumption made in [11] on the Kunneth type of differential forms.
@@
    Siegel operator for meromorphic modular forms (§5.3)
@@
The motivation of this paper comes from the pioneering work of Collino and Pirolla ([8], [9]). In [9], they discovered that the infinitesimal invariant of the Ceresa cycle in genus $3$ is essentially the equation of the universal plane quartic. Much later, the latter was studied independently from the viewpoint of modular forms ([7]); the connection was recently clarified in [18]. In the subsequent paper [8], Collino constructed higher Chow cycles in $g=2$ and $g=1$ by applying the K-theory elevator to the Ceresa cycle in $g=3$ successively, and expressed their infinitesimal invariants in terms of invariants of six points on ${{\mathbb{P}}}^{1}$ in the case $g=2$, and theta series in the case $g=1$ respectively. It was observed, again later and independently, that invariants of binary sextics are vector-valued Siegel modular forms ([6]). In a sense, like the age of Lemniscate, algebraic cycles have been ahead of modular forms. The present paper grew out of an effort to understand the idea of Collino from the viewpoint of modular forms.
I would like to thank Richard Hain for fruitful communication.
@@ 3. Higher Chow cycles
In this section we recall higher Chow cycles and normal functions. Our main references are [22] and [14]. Note that we would not need to define higher Chow groups, and hence avoid technicalities regarding the moving lemma in our presentation. (The exception is §4.3 and §4.4, but there higher Chow groups play only auxiliary roles.) We use the cubical presentation of higher Chow cycles.
@@ 3.1. Higher Chow cycles
For nullhomologous $Z$, the regulator $\nu(Z)\in J^{p,n}(X)$ is also called the Abel-Jacobi invariant of $Z$. This is a generalization of the classical Abel-Jacobi invariant in the case $n=0$. An integral formula in the case $n=1$ has been classically known (going back to Bloch); a generalization to the case $n\geq 2$ was given by Kerr, Lewis and Müller-Stach [22]. In this formula, a further proper intersection condition is imposed on $Z$ with regard to the real faces of ${\square}^{n}$ (see [22] §5.4). In §6 and later, where we implicitly use this formula, we tacitly assume that our cycles satisfy this condition.
@@ 4. The main construction
In this section we present and prove the full version of Theorem 1.1. In §4.1 we calculate the relevant Koszul complex as a complex of automorphic vector bundles. Our main construction is given in §4.2. The §4.3 and §4.4 are complements.
@@ 4.2. The main construction
Let us comment on some other possible constructions.
@@
When $g\geq 2$, we could also look at the non-primitive part of $\delta\nu_{Z}$. By (4.5), this produces Siegel modular forms of (virtual) weight $\wedge^{2}{{\rm Sym}}^{2}\otimes\det^{-1}$ when $g=3$, and weight ${{\rm Sym}}^{2}$ when $g=2$. The significance of considering the primitive part will be explained in §4.3 and §4.4.
@@
When the etale cover $S$ itself has a nice moduli interpretation and a satisfactory theory of “modular forms” on $S$ is available, we could directly study $\delta\nu_{Z}$ itself as such a generalized modular form on $S$ without pushing it down to ${\mathcal{A}_{\Gamma}}$. This is not the approach taken in this paper, but what we have in mind is the moduli space $\mathcal{M}_{3}\to\mathcal{A}_{3}$ of genus $3$ curves with the theory of Teichmüller modular forms ([21], [7]). See [18] for this direction.
@@
Another interesting source of normal functions is the extension classes of the mixed Hodge structures on the fundamental groups of curves ([16]). It is known that the first nontrivial extension class corresponds to the Ceresa cycle in the case $g=3$ ([29]) and to the Collino cycle in the case $g=2$ ([10]). The higher extension classes also give Siegel modular forms; the admissibility in this case is proved in [17].
4.3. On the primitive part
In this subsection we let $g=2,3$. Since Griffiths, the primitive part of normal functions has been used to detect various nontriviality of cycles. In our situation, this is expressed in the following form.
Proposition 4.9.
Let $Z$ be as in §4.2. Assume that for very general $s\in S$, the cycle $Z_{s}$ is algebraically equivalent to zero in the case $g=3$, and decomposable in the case $g=2$. Then $\delta\nu_{Z}^{+}\equiv 0$.
Here “very general points” means points in the complement of countably many divisors. A $(2,1)$-cycle on an abelian surface $A$ is called decomposable if it is in the image of the intersection product ${\rm Pic}(A)\otimes_{{\mathbb{Z}}}{\mathbb{C}}^{\ast}\to{\rm CH}^{2}(A,1)$.
Proof.
We first consider the case $g=3$. For a polarized abelian $3$-fold $A=X_{s}$, let $J^{2,0}(A)_{alg}$ be the sub abelian variety of the intermediate Jacobian $J^{2,0}(A)$ (usually written as $J^{3}(A)$) corresponding to the maximal sub ${\mathbb{Q}}$-Hodge structure of $H^{3}(A,{\mathbb{Q}})$ contained in $F^{1}H^{3}(A)$. The Abel-Jacobi invariant of a $1$-cycle algebraically equivalent to zero takes values in $J^{2,0}(A)_{alg}$. On the other hand, we let $J^{2,0}(A)_{Lef}$ be the sub abelian variety of $J^{2,0}(A)$ corresponding to the image of the Lefschetz operator $H^{1}(A,{\mathbb{Q}})\hookrightarrow H^{3}(A,{\mathbb{Q}})$. This is isogenous to ${\rm Pic}^{0}(A)$. Clearly we have $J^{2,0}(A)_{Lef}\subset J^{2,0}(A)_{alg}$, and $J^{2,0}(A)_{Lef}=J^{2,0}(A)_{alg}$ holds when $A$ is very general.
Now, by our assumption on $Z$, $\nu(Z_{s})$ takes values in $J^{2,0}(A)_{alg}=J^{2,0}(A)_{Lef}$ for very general $s\in S$. Since $J^{2,0}(A)_{Lef}$ is stable under deformation (while $J^{2,0}(A)_{alg}$ is not so), we see that $\nu(Z_{s})$ takes values in $J^{2,0}(A)_{Lef}$ for every $s\in S$. Since $\delta\nu_{Z}^{+}$ is the infinitesimal invariant of the image of $\nu_{Z}$ in the quotient of $J^{2,0}(X/S)$ by $J^{2,0}(X/S)_{Lef}$, it vanishes identically.
The case $g=2$ is similar: instead of $J^{2,0}(A)_{Lef}\subset J^{2,0}(A)_{alg}$ we use the sub tori ${\mathbb{Z}}H\otimes{\mathbb{C}}^{\ast}\subset{\rm NS}(A)\otimes{\mathbb{C}}^{\ast}$ of $J^{2,1}(A)$ where $H$ is the polarization. The Abel-Jacobi invariant of a decomposable cycle is contained in ${\rm NS}(A)\otimes{\mathbb{C}}^{\ast}$, while we have ${\rm NS}(A)={\mathbb{Z}}H$ for very general $A$. With this modification, the above argument works in the case $g=2$. ∎
Proposition 4.9 shows that passing to the primitive part of $\delta\nu_{Z}$ has the effect of dividing out the contribution from families of relatively accessible cycles as above.
4.4. Rigidity
In this subsection we let $1\leq g\leq 3$. We fix ${\mathcal{A}_{\Gamma}}$, but allow $S\to U$ to vary. The second significance of considering the primitive part is the following rigidity property.
Proposition 4.10.
There are only countably many Siegel modular forms that can be obtained as $f_{Z}^{(i)}$ for a cycle family $Z$ as in §4.2.
This is a consequence of some rigidity results in the theory of algebraic cycles, rather than a general property of normal functions (not necessarily of geometric origin). We begin with some preliminaries. For a polarized abelian $g$-fold $A$, we denote
$J^{2,3-g}(A)_{{\rm prim}}=\begin{cases}J^{2,0}(A)/J^{2,0}(A)_{Lef}&g=3,\\ J^{2,1}(A)/{\mathbb{Z}}H\otimes{\mathbb{C}}^{\ast}&g=2,\\ J^{2,2}(A)&g=1,\end{cases}$
where $J^{2,0}(A)_{Lef}$ and $H$ are as in the proof of Proposition 4.9. For a cycle family $Z$, we denote by $\nu_{Z}^{+}$ the image of $\nu_{Z}$ by the projection $J^{2,3-g}(X/S)\to J^{2,3-g}(X/S)_{{\rm prim}}$. The primitive part $\delta\nu_{Z}^{+}$ of $\delta\nu_{Z}$ is the infinitesimal invariant of $\nu_{Z}^{+}$.
Let $V$ be a small open set of $\mathcal{A}_{{\Gamma}}^{\circ}$ (in the classical topology). By a cycle family over $V$, we mean the restriction of a cycle family $Z$ on an etale base-changed family $X\to S$ to an open set $\tilde{V}\subset S$ which is projected isomorphically to $V$. We let $\mathcal{X}_{V}=\mathcal{X}_{{\mathcal{D}}}|_{V}$ be the universal family over $V$. Then, via the isomorphism $\tilde{V}\simeq V$, $\nu_{Z}^{+}|_{\tilde{V}}$ is regarded as a section of $J^{2,3-g}(\mathcal{X}_{V}/V)_{{\rm prim}}$ over $V$.
Lemma 4.11.
There are only countably many sections of $J^{2,3-g}(\mathcal{X}_{V}/V)_{{\rm prim}}$ that arises as $\nu_{Z}^{+}|_{\tilde{V}}$ for a cycle family $Z$ over $V$.
Proposition 4.10 follows from Lemma 4.11 because $f_{Z}^{(i)}$ is determined by its restriction to $V$, which in turn is a combination of $\delta\nu_{Z}^{+}|_{\tilde{V}}$ for various $\tilde{V}$ over $V$, hence a combination of some sections from a pool of countably many ones.
(Proof of Lemma 4.11).
We first consider the case $g=1$ which is simplest. The Beilinson rigidity ([2], see also [25] Corollary 3.3) says that the image of the Abel-Jacobi map ${\rm CH}^{2}(A,2)\to J^{2,2}(A)$ for an elliptic curve $A$ is countable. Hence the normal functions $\nu_{Z}|_{V}$ can take only countably many values at each point of $V$. Since a normal function is determined by its values at countably many (Zariski dense) points, this shows that there are only countably many sections of the form $\nu_{Z}|_{V}$.
Next let $g=3$. For a polarized abelian $3$-fold $A$, we consider the commutative diagram
[…]
where $J^{2,0}(A)_{tr}=J^{2,0}(A)/J^{2,0}(A)_{alg}$, ${\rm CH}^{2}(A)_{{\rm hom}}$ is the Chow group of nullhomologous $1$-cycles, and ${\rm Griff}^{2}(A)$ is their Griffiths group. If $A$ is very general, we have $J^{2,0}(A)_{{\rm prim}}=J^{2,0}(A)_{tr}$. Since ${\rm Griff}^{2}(A)$ is a countable group, the image of the primitive Abel-Jacobi map $\nu^{+}$ is countable for very general $A$. As in the case $g=1$, this implies that there are at most countably many sections of the form $\nu^{+}_{Z}|_{V}$.
The case $g=2$ is similar. We replace the above diagram by
[…]
where ${\rm CH}^{2}(A,1)_{{\rm ind}}={\rm CH}^{2}(A,1)/{\rm Pic}(A)\otimes{\mathbb{C}}^{\ast}$ is the indecomposable part of ${\rm CH}^{2}(A,1)$ and $J^{2,1}(A)_{tr}=J^{2,1}(A)/{\rm NS}(A)\otimes{\mathbb{C}}^{\ast}$. Again we have $J^{2,1}(A)_{tr}=J^{2,1}(A)_{{\rm prim}}$ for very general $A$, and now the image of $\nu_{tr}$ is countable by [25] Corollary 3.6. ∎
@@ 6.1. Rank $1$ degeneration
We will come back to (1) – (5) from a general point of view on another occasion. For the moment, these are taken as assumptions in the rest of this paper, which are satisfied for the examples in [19].
@@ 6.3. The statement
Recall from Theorem 4.2 that the cycle family $Z$ gives a meromorphic Siegel modular form $f_{Z}^{(i)}$ of weight ${{\rm Sym}}^{4i}\otimes{\rm det^{-i}}$ on ${\mathcal{A}_{\Gamma}}$ for each $1\leq i\leq d$. In Lemma 8.1, we will prove that $f_{Z}^{(i)}$ is holomorphic at a general point of $\Delta_{I}$. However, as explained in §5.3, this is still weaker than the regularity at $I$ in the sense of Definition 5.4. Let us assume that the divisor components of ${\mathcal{A}_{\Gamma}}-U$ do not contain ${\mathcal{A}_{I}}$ in their closure. Then, by Lemma 5.3 (1), $f_{Z}^{(i)}$ is regular at $I$. This enables us to apply the meromorphic Siegel operator to $f_{Z}^{(i)}$. By Proposition 5.6, the resulting $\Phi_{I}f_{Z}^{(i)}$ is a meromorphic Siegel modular form over ${\mathcal{A}_{I}}$, again of weight ${{\rm Sym}}^{4i}\otimes{\rm det}^{-i}$.
On the other hand, if consider the cycle family $Z_{I}$ on $X_{I}\to S_{I}$ and take the same product construction as $\delta\nu_{Z}^{+}\rightsquigarrow f_{Z}^{(i)}$ for the etale cover $S_{I}\to{\widetilde{\mathcal{A}}_{I}}$ of degree $d$ (rather than $S_{I}\to{\mathcal{A}_{I}}$ which in general has larger degree), we obtain a meromorphic section $f_{Z_{I}}^{(i)}$ of $\phi_{0}^{\ast}({{\rm Sym}}^{4i}{\mathcal{E}}_{I}\otimes{\mathcal{L}}_{I}^{\otimes-i})$ over ${\widetilde{\mathcal{A}}_{I}}$, where ${\mathcal{E}}_{I}$ and ${\mathcal{L}}_{I}$ are the Hodge bundles on ${\mathcal{A}_{I}}$.
@@ 8.4. The case $g=3$
The proof of Theorem 6.2 in the case $g=3$ is similar to the case $g=2$. In this case $(H^{1}\mathcal{K}_{3})^{I}$ no longer coincides with $(\mathcal{H}^{1,2}_{{\rm prim}})^{I}\otimes\Omega^{1}(\log\Delta_{I})^{I}$, but is still realized as its sub vector bundle by Proposition 7.4. The cusp ${\mathcal{D}}_{I}$ is now the $3$-dimensional Siegel upper half space, so we use $3$-variable coordinates $(\tau_{ij})_{i,j}$ instead of a single variable $\tau$. Then we replace $\partial_{\tau}\otimes\omega$ by $\partial_{\tau_{ij}}\otimes\omega$ as the test dual sections where now $\omega_{t}\in H^{2,1}_{{\rm prim}}(X_{t})$ converging to $(2\pi i)^{-1}\eta_{0}\wedge(dz/z)$ with $\eta_{0}\in H^{1,1}_{{\rm prim}}(B)$. After this modification, the proof of Proposition 8.4 is valid in the case $g=3$. The step deducing Theorem 6.2 from Proposition 8.4 remains the same.